A rocket is headed away from Earth at a speed of 0.80c. The rocket fires a small

payload at a speed of 0.70c (relative to the rocket) aimed away from Earth. How

fast is the payload moving relative to Earth?

(a) 1.50c;

(b) a little less than 1.50c;

(c) a little over c;

(d) a little under c;

(e) 0.75c.


1
Expert's answer
2021-06-11T11:29:40-0400

"v = 0.8c \space\space and \space\space u' =0.7c"

"v, u' =>" along the same direction


"u = \large\frac{v + u'}{1 + \large\frac{vu'}{c^2}}= \frac{1.5c}{1+0.56} =" "0.96c"

(d) a little under c